OpenStudy (anonymous):

can someone explain how: m(2d)^2 =4md^2

OpenStudy (anonymous):

Because you need to square everything inside the ( )

OpenStudy (anonymous):

Because the ^2 isn't just positioned next to "d" but the ( ) surrounding 2d

OpenStudy (anonymous):

2^2 = 4 d^2 = d^2 then you get 4md^2...and where you place the 4 doesn't matter as everything is being multiplied..get it?

OpenStudy (anonymous):

ohhh i get it! it only squares the 2!

OpenStudy (baru):

what jaeuni means is \(m(2d)^2=m\times 2^2 \times d^2\)

OpenStudy (anonymous):

no it squares everything inside of ( )...so you're squaring both the 2 and d which is why the d also becomes d^2

OpenStudy (anonymous):

yes thats what i meant, it also squares the d,. Thanks!

OpenStudy (anonymous):

np :)